If one of the lines given by 6x2-xy+4cy2=0 is 3x+4y=0,then c=

Grade:8

2 Answers

Arun
3 years ago
If 3x +4y = 0 is one line
 
3x = -4y
 
Now put this in Equation of pair of line
 
96y²/9  + 4y²/3 + 4cy² = 0
 
96 + 12 + 36c = 0
 
36c = -108
c = -3
 
Hope it helps
Vikas TU
3 years ago
Dear student 
6x^2 -xy +4cy^2 =0
put y = -(3x/4)
then 6x^2 +3x^2/4 +4c*(9x^2/16)=0
c = -3 
Good Luck 
Cheers